OpenStudy (anonymous):

an=2+(n-1)6 i need to know what the an means and a1 means i am doing arithmetic sequence and geometric sequence

jimthompson5910 (jim_thompson5910):

a1 is the first term an is the nth term

jimthompson5910 (jim_thompson5910):

the general arithmetic sequence is given by an = a1 + (n-1)*d so we see from an=2+(n-1)6 that a1 = 2 and d =6

OpenStudy (anonymous):

yes so what does an mean and a1 mean and d means

jimthompson5910 (jim_thompson5910):

with an=2+(n-1)6, the first term is 2, and the common difference is 6

jimthompson5910 (jim_thompson5910):

so the sequence is 2, 8, 14, 20, ...

jimthompson5910 (jim_thompson5910):

an is just the general term (or nth term)

OpenStudy (anonymous):

ok so what is the d

jimthompson5910 (jim_thompson5910):

d = 6

jimthompson5910 (jim_thompson5910):

that's the common difference

jimthompson5910 (jim_thompson5910):

ie what you add/subtract to each term

OpenStudy (anonymous):

so if i had to write an arithmetic sequence with 5 terms what would it look like

OpenStudy (anonymous):

jim you still here

jimthompson5910 (jim_thompson5910):

If you want to write the first 5 terms of an=2+(n-1)6, start with 2, and add 6 to each current term to get the next term

jimthompson5910 (jim_thompson5910):

2+6 = 8 8+6 = 14 14+6 = 20 20+6 = 26

jimthompson5910 (jim_thompson5910):

so the first five terms of an=2+(n-1)6 are 2, 8, 14, 20, 26

OpenStudy (anonymous):

so the problem alone would be the sequence

jimthompson5910 (jim_thompson5910):

what do you mean

OpenStudy (anonymous):

i am needing an arithmetic sequence using 5 numbers (terms)

jimthompson5910 (jim_thompson5910):

i just gave you one 2, 8, 14, 20, 26

OpenStudy (anonymous):

ok sorry new to algebra just started this year thanks for all your help

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

is a geometric sequence the same as the arithmetic sequence

jimthompson5910 (jim_thompson5910):

no, a geometric sequence is generated by multiplying (instead of adding) some constant number to each term

jimthompson5910 (jim_thompson5910):

an example geometric sequence is 1, 2, 4, 8, 16 because we're multiplying each term by 2 to get the next term

OpenStudy (anonymous):

an=5*3n-1

OpenStudy (anonymous):

so we would be mutiplying each term by two in this one right

jimthompson5910 (jim_thompson5910):

in the case of an=5*3^(n-1), you start with 5, and you multiply it by 3 to get 15, then you multiply that by 3 to get 45, etc etc to get this sequence 5, 15, 45, 135, 405

OpenStudy (anonymous):

so what would the n be

jimthompson5910 (jim_thompson5910):

n is a variable and it represents the current term for instance, for the first term, n = 1 for the second term, n = 2, for the third term, n = 3, etc

OpenStudy (anonymous):

an is the nth term in this one correct and a1 is the first term right

jimthompson5910 (jim_thompson5910):

an is the nth term a1 or just 'a' is the first term

OpenStudy (anonymous):

ok i think i am getting the hang of this

jimthompson5910 (jim_thompson5910):

ok great

OpenStudy (anonymous):

with an equation an =3*5n-1 what would the sequence be. multiply the first by five and so on and so forth right

OpenStudy (anonymous):

i came up with 3,15,75,425,2125 is that correct

OpenStudy (anonymous):

5 would be the common ratio i think

OpenStudy (anonymous):

is that correct jim

jimthompson5910 (jim_thompson5910):

for an =3*5n-1, the first five terms are 3, 15, 75, 375, 1875

OpenStudy (anonymous):

dont you multiply everything by 5 except the first one right

jimthompson5910 (jim_thompson5910):

no, you multiply 3 by 5 to get 15 15 by 5 to get 75 75 by 5 to get 375 etc

OpenStudy (anonymous):

yeah thats what i thought just messed up the math on 75*5 thats where i messed up

jimthompson5910 (jim_thompson5910):

i gotcha

OpenStudy (anonymous):

thanks for all your help jim i think i understand this now

jimthompson5910 (jim_thompson5910):

sure thing, glad it's all clicking now